Quantitative Problem: Assume that interest rate parity holds. In the spot market 1 Japanese yen = $0.009, while in the 180-day forward market 1 Japanese yen = $0.0096. 180-day risk-free securities yield 1.50% in Japan.

What is the yield on 180-day risk-free securities in the United States? Do not round intermediate calculations. Round your answer to two decimal places.

%?
Answer #1

(1 + domestic risk free rate) = (1 + foreign risk free rate) * [spot exchange rate / forward exchange rate]

In this question, treat Japan as the domestic country (because exchange rate is quoted in terms of U.S. dollars per Japanese Yen). Applying the interest rate parity, we have:

(1 + 0.015) = (1 + U.S. risk-free rate) * [0.009 / 0.0096]

1.015 = (1 + U.S. risk-free rate) * 0.9375

1 + U.S. risk-free rate = 1.015 / 0.9375

U.S. risk-free rate = 1.0827 - 1 = 0.0827, or 8.27%